# Attribute or Abstain: Large Language Models as Long Document Assistants

### v1

@@ 4.1 RQ1: What are optimal approaches to attribution in long document tasks?
Flan-T5-XL has higher average scores than GPT-3.5 and GPT-4, while the Longchat scores are lower. For GPT-3.5, GPT-4 and Flan-T5-XL, citation/reduced-citation results in the best evidence quality on average and most datasets, and retrieve-then-read performs worst. Post-hoc works best for GovReport and Longchat.
@@
Comparing reduced-post-hoc to post-hoc and reduced-citation to citation, we find that response quality is mostly better for the non-reduced variant. Regarding evidence quality, our findings are not as clear, as reduced-citation results in the best average evidence quality for GPT-3.5.
@@ Discussion
Citation or reduced-citation result in the best average evidence quality, while not hurting response quality, in line with recent work showing LLM capabilities for retrieval Ma et al. (2023). The GovReport task and the Longchat model are exceptions to this, as post-hoc results in better evidence quality in these cases. For GovReport, the higher evidence quality with post-hoc can be explained with the "repetitive" nature of the summarization task, since the high overlap between response statements and document provides good conditions for retrievers to find evidence. For small models such as Longchat, related work has shown that they lack instruction following capability to perform evidence extraction Gao et al. (2023b); Schimanski et al. (2024), making post-hoc the better approach for the model.
@@
Similar to Xu et al. (2024) and Bai et al. (2023), we have not found a general beneficial effect of context reduction on response quality. The positive results from Agrawal et al. (2024) were obtained in a multi-document setting, where there is no logical coherence between the input documents. In contrast, the logical coherence in long documents can be disrupted through reduction, which can make processing of the reduced document more difficult.
@@ 4.2 RQ2: Do LLMs exhibit positional biases in attribution?
Figure 3 (top) shows the predicted evidence distributions from citation [^12] We focus on citation because only in this approach the LLM performs evidence retrieval. and the gold distributions. [^13] For GovReport, we matched summary sentences to paragraphs using BM25 to obtain gold evidence, see §3.1 The models generally follow the gold evidence distribution, with Longchat showing the strongest deviation. GovReport is an exception: all models show a higher focus on the beginning of the document, especially Flan-T5-XL.
@@ 4.3 RQ3: What is the relation between evidence quality and response quality?
citation 0 0 14 17 14 4 24 -1
GPT-4 post-hoc 4 1 24 26 9 11 16 0
citation 3 5 9 20 11 3 18 1
Flan-T5 post-hoc -1 2 14 13 8 0 8 0
citation 2 1 11 11 11 4 9 -1
@@
Attributability evaluation models Honovich et al. (2022); Yue et al. (2023); Tang et al. (2024) can evaluate evidence quality without external reference. If evidence quality were positively correlated with response quality, this could be used to abstain from low-quality responses. We test this with selective prediction El-Yaniv and Wiener (2010), using attributability scores as an estimate of confidence.
@@
We begin by filtering responses that don’t require evidence (e.g. “unanswerable” on QASPER or “not mentioned” on ContractNLI), as attributability cannot be evaluated on these. We sort the remaining $l$ predictions by attributability, [^14] We used MinicheckTang et al. (2024) for QASPER and TRUE Honovich et al. (2022) for all other datasets (§3.2.1) obtaining the set of responses $\mathcal{R}_{sel}$ in descending order. For each $t\in\{0...l\}$, we compute average response quality [^15] As filtering responses that do not require evidence produces a strong class imbalance, we use micro F1 instead of macro F1 for Wice and ContractNLI on the subset of $\mathcal{R}_{sel}$ up to $t$ and the coverage (the proportion of responses evaluated). We evalate the confidence estimation by the area under the response quality-coverage curve (AUC) Chen et al. (2023).
@@
Table 4 shows the AUC difference between ordering predictions by attributability and a random order baseline (mean of 10 repeats). We see that attributability is an effective estimate of confidence on Natural Questions, Evidence Inference and ContractNLI, and, to some extent on Wice. On QASPER, and GovReport, the difference to the random baseline is small.
@@
For instances annotated as “unanswerable”, “not supported” or “not mentioned”, there is no annotated evidence. If models give different responses, the evidence quality should be low, and these should be filtered in selective prediction. Fig. 4 shows the distribution of gold responses in selective prediction. The proportion of such instances without sufficient evidence decreases with lower coverage for Natural Questions, Wice and ContractNLI but, surprisingly, not for QASPER.
@@ Discussion
We explain the differences in the relationship between evidence quality and response quality by the varying dataset complexity. While Natural Questions, ContractNLI and Evidence Inference focus on single facts (e.g. a single entity or a specific contractual obligation), Wice and QASPER instances often contain multiple facts (e.g. an enumeration or multiple subclaims), which is also reflected in the number of evidence segments per instance (Table 1). Models respond correctly, but fail to point to all necessary evidence. This is in line with related work on attribution: While Bohnet et al. (2022) found that response quality can be improved by attributability-based reranking on Natural Questions, Gao et al. (2023b) did not find this on their more complex benchmark.
@@ 6 Limitations
While our attributability model selection experiments showed that these models obtain good accuracy, our analysis in 4.3 showed that edge cases are not yet handled well. Research into solving such edge cases is a promising direction for future work.
@@
We compiled a benchmark with diverse tasks, domains, response and document lengths, but naturally, we were not able to cover all variations of these properties. Many long document datasets and benchmarks are available (e.g. Li et al. 2023b; An et al. 2023; Shaham et al. 2023), but only few contain annotated evidence, which we required for the positional bias analysis in our paper (not finding a long document summarization dataset with annotated evidence, we resorted to GovReport). Given that we found only limited positional biases, extending our work to more datasets with and without annotated evidence is an interesting direction for future work.
@@ A.1 Correlation between Evidence Position and Response Quality
citation -0.26 -0.28 -0.42 -0.30 -0.54 -0.81*
Table 5: Correlation between response quality and position of annotated evidence. See §4.2 for details.
@@ A.2 System-level Correlation of Attributability and Response Quality
QSP $0.86$ ($6.8\times 10^{-3}$)
NQ $0.99$ ($2.7\times 10^{-6}$)
EI $0.97$ ($3.9\times 10^{-5}$)
WIC $0.90$ ($2.3\times 10^{-3}$)
CNLI $0.93$ ($9.4\times 10^{-4}$)
GR $0.18$ ($6.7\times 10^{-1}$)
Avg $0.98$ ($1.2\times 10^{-5}$)
@@ D.3 Annotation Instructions
As mentioned in 3.2.1, we manually annotated predictions on QASPER, Natural Questions and GovReport. The annotation instructions are shown in Fig. 5.
@@ Appendix G Technical Details
Gemma 22 22 0 20 1 12.51
GritLM 26 46 0 22 0 17.44
Mistral 31 47 0 0.25 0 19.33
Vicuna-7B 21 42 0 23 2 16.67
Flan-T5-XXL 26 28 36 12 1 18.32
LongChat 21 52 0 0.25 16 22.31
Llama3 17 60 0 23 0 18.78
Table 18: results of OS Prompt Model Selection

### v2

@@ 5.1 RQ1: What are optimal approaches to attribution in long document tasks?
Flan-T5-XL has higher average scores than GPT-3.5 and GPT-4, while the Longchat and Mistral scores are lower. For GPT-3.5, GPT-4 and Flan-T5, citation / reduced-citation results in the best evidence quality on average and most datasets, and retrieve-then-read performs worst. Post-hoc works best for Longchat and Mistral, and for all models on GovReport.
@@
Does performance depend on document length?
Tables 7 and 8 show the correlation between response quality or evidence quality and document length, which is mostly negative. A notable exception is GPT-4 evaluated for response quality, where correlation with document length is positive on all tasks but GovReport.
@@
Comparing reduced-post-hoc to post-hoc and reduced-citation to citation, we find that response quality is mostly better for the non-reduced variant, with the exception of Mistral. Regarding evidence quality, reduction only helps for GPT-3.5-citation, Longchat-citation and -post-hoc and Mistral-citation.
@@ Discussion
Citation or reduced-citation result in the best average evidence quality, while not hurting response quality, in line with recent work showing LLM capabilities for retrieval Ma et al. (2023). The GovReport task and the small models Longchat and Mistral are exceptions to this, as post-hoc results in better evidence quality in these cases. For GovReport, the higher evidence quality with post-hoc can be explained with the "repetitive" nature of the summarization task, since the high overlap between response statements and document provides good conditions for retrievers to find evidence. For the small models, related work has shown that they lack instruction following capability to perform evidence extraction Gao et al. (2023b); Schimanski et al. (2024), making post-hoc the better approach for the model.
@@
Regarding the relation between input length and performance, we mostly found negative correlation, similar to Bai et al. (2023) and Kwan et al. (2023). Still, similar to Xu et al. (2024) and Bai et al. (2023), we have not found a general beneficial effect of input reduction on performance. The positive results from Agrawal et al. (2024) were obtained in a multi-document setting, where there is no logical coherence between the input documents. In contrast, the logical coherence in long documents can be disrupted through reduction, which can make processing of the reduced document more difficult.
@@ 5.2 RQ2: Do LLMs exhibit positional biases in attribution?
Figure 3 (top) shows the predicted evidence distributions from citation [^14] We focus on citation because only in this approach the LLM performs evidence retrieval. and the gold distributions. [^15] For GovReport, we matched summary sentences to paragraphs using BM25 to obtain gold evidence, see §4.1 The models generally follow the gold evidence distribution, with Longchat showing the strongest deviation. GovReport is an exception: all models show a higher focus on the beginning of the document, especially Flan-T5-XL. Figures 5 and 6 show the relation of evidence quality and the position of gold or predicted evidence, respectively. Neither of them show a clear dependence of evidence quality on evidence position.
@@ 5.3 RQ3: What is the relation between evidence quality and response quality?
citation 2 1 14 18 13 4 25 -1
GPT-4 post-hoc 4 1 20 27 7 10 16 0
citation 4 3 6 18 10 7 20 1
Flan-T5 post-hoc 0 2 13 12 7 1 8 0
citation 2 1 11 10 11 3 8 -1
@@
Attributability evaluation models Honovich et al. (2022); Yue et al. (2023); Tang et al. (2024) can evaluate evidence quality without external reference. If evidence quality were positively correlated with response quality, this could be used to abstain from low-quality responses. We test this with selective prediction El-Yaniv and Wiener (2010), using evidence quality scores from attributability evaluation as an estimate of confidence.
@@
We begin by filtering predicted responses that don’t require evidence161616“Unanswerable” on QASPER and Natural Questions, “not supported” on Wice, “not mentioned” on ContractNLI., as attributability cannot be evaluated on these. We sort the remaining $l$ predictions by evidence quality [^17] We evaluated evidence quality as attributability for all datasets, as it does not require annotated evidence. We used MinicheckTang et al. (2024) for QASPER and TRUE Honovich et al. (2022) for all other datasets (§4.2.3) in descending order, obtaining the ordered set of responses $\mathcal{R}_{sel}$. For each $t\in\{0...l\}$, we compute average response quality [^18] As filtering responses that do not require evidence produces a strong class imbalance, we use micro F1 instead of macro F1 for Wice and ContractNLI on the subset of $\mathcal{R}_{sel}$ up to $t$ and the coverage (the proportion of responses evaluated). We evaluate the estimation of confidence through evidence quality by computing the area under the response quality-coverage curve (AUC) Chen et al. (2023).
@@
Table 4 shows the AUC difference between ordering predictions by attributability and a random order baseline (mean of 10 repeats). We see that attributability is an effective estimate of confidence on Natural Questions, Evidence Inference and ContractNLI, and to some extent on Wice. On QASPER, and GovReport, the difference to the random baseline is small.
@@
For instances annotated as “unanswerable”, “not supported” or “not mentioned”, there is no annotated evidence. If models give different responses, the evidence quality should be low, and these should be filtered in selective prediction. Fig. 4 shows the distribution of gold responses in selective prediction. As expected, the proportion of such instances without sufficient evidence decreases with lower coverage for Natural Questions, Wice and ContractNLI but, surprisingly, not for QASPER.
@@ Discussion
We explain the differences in the relationship between evidence quality and response quality by the varying dataset complexity. While Natural Questions, ContractNLI and Evidence Inference focus on single facts (e.g. a single entity or a specific contractual obligation), Wice and QASPER instances often contain multiple facts (e.g. an enumeration or multiple subclaims), which is also reflected in the higher number of evidence segments per instance (Table 1). Models respond correctly, but fail to point to all necessary evidence. This is in line with related work on attribution: While Bohnet et al. (2022) found that response quality can be improved by attributability-based reranking on Natural Questions, Gao et al. (2023b) did not find this on their more complex benchmark.
@@ 7 Limitations
While our attributability model selection experiments showed that these models obtain good accuracy, our analysis in 5.3 showed that edge cases are not yet handled well. Research into solving such edge cases is a promising direction for future work.
@@
We compiled a benchmark with diverse tasks, domains, response and document lengths, but naturally, we were not able to cover all variations of these properties. Many long document datasets and benchmarks are available (e.g. Li et al. 2023b; An et al. 2024; Shaham et al. 2023), but only few contain annotated evidence, which we required for the positional bias analysis in our paper (not finding a long document summarization dataset with annotated evidence, we resorted to GovReport). Given that we found only limited positional biases, extending our work to more datasets with and without annotated evidence is an interesting direction for future work.
@@ A.1 Correlation between Evidence Position and Response Quality
citation -0.49 -0.28 -0.42 -0.30 -0.54 -0.61
Mistral post-hoc -0.53 0.39 -0.59 0.04 0.09 -0.92*
citation -0.35 0.31 -0.37 -0.58 0.47 -0.82*
Table 5: Pearson correlation between response quality and position of annotated evidence. * indicates significance ($p<0.05$). See §5.2 for details.
@@ A.2 System-level Correlation of Attributability and Response Quality
QSP $0.93$ ($7.7\times 10^{-5}$)
NQ $0.99$ ($3.6\times 10^{-8}$)
EI $0.93$ ($7.9\times 10^{-5}$)
WIC $0.96$ ($1.2\times 10^{-5}$)
CNLI $0.93$ ($1.2\times 10^{-4}$)
GR $0.36$ ($3.0\times 10^{-1}$)
mean $0.98$ ($1.3\times 10^{-6}$)
@@
A.3 Overlap between citation and post-hoc evidence
A.4 Relation between Performance and Input Document Length
See Tables 7 and 8. See §5.1 for analysis and discussion.
QSP NQ EI WIC CNLI GR
GPT-3.5 citation 0.04 -0.01 -0.08* -0.04 0.05* -0.08*
post-hoc 0.08* -0.00 -0.05 -0.05 0.05* -0.04
GPT-4 citation 0.29* 0.11 0.13 0.08 0.34* -0.34*
post-hoc 0.28* 0.17 0.13 0.11 0.39* -0.26*
Flan-T5 citation -0.02 -0.10* -0.07* 0.05 0.00 -0.07*
post-hoc -0.03 -0.10* -0.05 -0.08 -0.02 -0.11*
Longchat citation -0.05 -0.05* -0.04 0.01 0.04 -0.15*
post-hoc -0.02 -0.13* -0.04 0.03 -0.06* -0.13*
Mistral citation -0.03 -0.07* -0.07* -0.07 0.00 -0.12*
post-hoc -0.01 -0.06* -0.07* -0.04 -0.02 -0.06
Table 7: Pearson correlation between response quality and document length. * indicates significance ($p<0.05$). See §5.1 for analysis and discussion.
QSP NQ EI WIC CNLI GR
GPT-3.5 citation -0.08* -0.13* -0.10* -0.20* -0.04 -0.24*
post-hoc -0.03 -0.10* -0.01 -0.11* -0.07* -0.23*
GPT-4 citation 0.04 -0.17 -0.09 0.09 0.15 -0.10
post-hoc -0.13 -0.05 -0.07 0.10 0.01 -0.12
Flan-T5 citation -0.04 -0.08 -0.10* -0.09 -0.07* -0.11*
post-hoc -0.06* -0.20* -0.00 -0.15* -0.13* -0.10*
Longchat citation -0.13* -0.09* -0.12* -0.05 0.08* -0.03
post-hoc -0.01 -0.13* 0.00 0.03 0.04* -0.18*
Mistral citation -0.18* -0.17* -0.13* -0.19* -0.09* -0.05
post-hoc -0.08* -0.10* -0.04 -0.16* -0.06* -0.10*
Table 8: Pearson correlation between evidence quality and document length. * indicates significance ($p<0.05$). See §5.1 for analysis and discussion.
A.5 Effect of Evidence Position on Evidence Quality
To analyze the effect of the position of gold or predicted evidence on evidence quality, we grouped predictions from models under the citation approach into 5 bins based on the relative position of gold or predicted evidence in the document. We then computed mean evidence quality separately for each bin. Figures 5 and 6 show the results. A “Lost in the Middle” effect is not visible.
Figure 5: Evidence quality by position of gold evidence in the document. “Lost in the Middle” effect is not visible. For more information, see §A.5.
Figure 6: Evidence quality by position of predicted evidence in the document. A “Lost in the Middle” effect is not visible. For more information, see §A.5.
@@ D.3 Annotation Instructions
As mentioned in 4.2.3, we manually annotated predictions on QASPER, Natural Questions and GovReport. The annotation instructions are shown in Fig. 7.
@@ Appendix G Technical Details
Detecting abstaining
Based on Slobodkin et al. (2023) and our own inspection of model responses, we compiled two sets of keyphrases to detect abstaining. If any of these key was found in a response, it was set to “unanswerable”. The first set of keyphrases is shown in Table 26, and was used as is. The second set of keyphrases was constructed by combining all verbs and adverbs in Table 27 as follows:
  1. 1.
    “not <verb>”
  2. 2.
    “not <adverb> <verb>”
Verbs were used in their base form and as participles (e.g. “mention” $\rightarrow$ “mentioned”). This means that each combination of verb and adverb results in four keyphrases. For example, the verb “provide” and the adverb “explicitly” resulted in “not provide”, “not explicitly provide”, “not provided”, “not explicitly provided”.
@@
Gemma-7b-it 22 2 49 18 0 17
GritLM-7B 26 6 48 20 0 18
Mistral-7B-Instruct-v0.2 31 32 48 24 0 25
Vicuna-7B-v1.5-16K 21 12 49 21 2 19
Flan-T5-XXL 26 19 66 12 1 22
LongChat-7B-v1.5-32K 21 18 48 22 13 24
Llama3-8B 17 17 47 22 0 19
Table 20: Results of model selection for prompted open source models.
@@
unanswerable
n/a
i don’t know
idk
not known
answer not in context
unknown
no answer
it is unknown
the answer is unknown
unavailable
not clear
i cannot provide
i cannot directly provide
i cannot answer
i cannot display
not clear
not available
not readily available
Table 26: Simple keyphrases to detect abstaining. See §G for details on how these were used.
Verbs Adverbs
provide explicitly
mention specifically
state directly
specify clearly
define
report
name
offer
Table 27: Verbs and adverbs that were combined to phrases to detect abstaining. See §G for details on how these were used.